The sole owner of Auto Moto Spares needs to increase production as a result of increasing demand from motor accessory shops. Current capacity has been at 4, 000 units per year. Each spare is sold to the retailers for N$60. Production costs are:Direct labour N$20 Direct materials N$16 Fixed costs N$58, 000 The owner is considering two options for expansion: Option 1 Extend the existing premises but keep the same method of production. This would increase fixed costs by N$30, 000 per year, but direct costs would remain unchanged. Capacity would double. Option 2 Purchase new machinery, which will speed up the production process and cut down on wasted materials. Fixed costs would rise by N$8, 000 per year, but direct costs would be reduced by N$4 per unit. Output capacity would increase by 60%. 3.1 Identify the Break-even point in units for each option. 8) 3.2 Which option would you advice the owner to choose? Motivate your answer. (4


1
Expert's answer
2021-05-03T10:49:40-0400

3.1 The Break-even point in units for option 1:

"BEP = \frac{88,000} {60 - (20 + 16)} = 3,666.67" units.

For option 2:

"BEP = \frac{66,000} {60 - (20 + 16 - 4)} = 2,357.14" units.

3.2 I would advice the owner to choose the option 2, because the fixed costs increase less, but direct costs decrease. It also has lower break-even point.